OpenStudy (anonymous):

Given the equation 2x +4 = 4x -2, select the reasoning that correctly solves for x. A.add 2, subtract 2x, then divide by 2 B.add 2, subtract 4x, then divide by -2 C.subtract 4, subtract 2x, then divide by -2 D.subtract 4, subtracts 4x, then divide by 2.

OpenStudy (anonymous):

A. Cause we're just solving for x, to isolate it. So 2x + 4 = 4x -2 You first subtract add 2 to both sides (whatever you do to one side you ave to do to another) 2x + 4 + 2 = 4x - 2 + 2 2x + 4 + 2 = 4x Then you subtract 2x from both sides so all your x coefficients are on the same side. 2x + 4 + 2 - 2x = 4x - 2x 4 + 2 = 4x - 2x Now, you combine your like terms. 6 = 2x Divide by -2 to completely solve for x. x = 3

OpenStudy (anonymous):

Ok thank you

OpenStudy (the_fizicx99):

You would divide by 2, not -2

OpenStudy (anonymous):

That's what I meant sorry. Lol no prob x

OpenStudy (anonymous):

Ok so I divide by 2 and should get -3

OpenStudy (the_fizicx99):

No, 6 = 2x /2 /2 x = 3

OpenStudy (the_fizicx99):

If it were negative it would've been x = -3

OpenStudy (anonymous):

I agree that the answer is A

OpenStudy (the_fizicx99):

Yesh the answer is a, though there was a typo :3

OpenStudy (anonymous):

Oh okay thanks

OpenStudy (anonymous):

-(3x-2)=-4 (x+1) -4 do I distribute?

OpenStudy (the_fizicx99):

Post it on a different question, its starts to lag after a while